A cruise boat travels 72 miles down stream in two hours and returns upstream in 6 hours. Find the rate of the stream.
D=RT so R=D/T 72/2=36 72/6=12 36-12=24 (speed of boat on still water?)
I get lost here.

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(48556)

A cruise boat travels 72 miles down stream in two hours and returns upstream in 6 hours. Find the rate of the stream.
-----------------------
Down stream DATA:
distance = 72 miles ; time = 2 hrs: rate = d/t = 72/2 = 36 mph
---------------------------
Up stream DATA:
distance = 72 miles ; time = 6 hrs. rate = 72/6 = 12 mph
------------------------
Equations
Down stream rate: b + c = 36 mph
Upstream rate...: b - c = 12
-------------------------------------
Add to get 2b = 48
b = 24 mph (rate of the boat in still water)
---
Substitute to find "c":
b+c = 24 + c = 36
c = 36-24 = 12 mph (speed of the current)
===============================================
Cheers,
Stan H.

Answer by Alan3354(21605)

A cruise boat travels 72 miles down stream in two hours and returns upstream in 6 hours. Find the rate of the stream.
D=RT so R=D/T 72/2=36 72/6=12 36-12=24 (speed of boat on still water?)
----------------
Going downstream, the speed of the stream is added to the speed of the boat.
Upstream, it's subtracted.
72 = (b+s)*2
72 = (b-s)*6
b+s = 36
b-s = 12
--------
2b = 48
b = 24 mph (speed of boat)
s = 12 mph (speed of stream)